Translate the following problems into algebraic equations :

1) A woman buys three pairs of jeans for $99.84, including tax. If each pair of jeans cost $32, what was the sales tax rate?
2) The final grade for a particular course is calculated by dividing total points earned by total points available. There are 120 points available in the first half of the course, and 140 points in the second half of the course. If a student earns 81% in the first half of the course, how many points must the student earn in the second half in order to earn an 80% B final grade?

Answer by stanbon(75887)

1) A woman buys three pairs of jeans for $99.84, including tax. If each pair of jeans cost $32, what was the sales tax rate?
-------------------
Cost of 3 jeans = 3*32 = 96
Amount of tax: 99.84-96 = 3.84
---------------------------------
% taz = 3.84/96 = 4%
========================

2) The final grade for a particular course is calculated by dividing total points earned by total points available. There are 120 points available in the first half of the course, and 140 points in the second half of the course. If a student earns 81% in the first half of the course, how many points must the student earn in the second half in order to earn an 80% B final grade?
---------------
final grade = (total pts)/(pts. available)
pts available = 120 + 140 = 260
--------------------------------
EQUATION:
1st half: 0.81 = total pts/120
total pts = 0.81*120 = 97.2 (pts. needed in 1st half to get 81%)
-----------------------------------
total pts needed for the course:
0.80 = total pts/260
total pts = 208 (pts. needed to get final grade of 80%)
---------------------------
pts. needed in the 2nd half = 208-97.2 = 110.8
================================================
Cheers,
Stan H.